# Quantum thermocouples: nonlocal conversion and control of heat in semiconductor nanostructures

### v1

@@ IV Interacting systems
In some quantum systems, single-particle interactions become important and cannot be treated in a mean-field level. This is the case of quantum dot systems weakly coupled to electronic reservoirs [233], where strong Coulomb interactions stabilize the number of allowed electrons (Coulomb blockade regime) up to single-charge fluctuations through the tunnel barriers. Then, Coulomb interactions between electrons in capacitively coupled conductors or interactions with bosonic excitations (phonons, photons, magnons) need to be considered with a single-particle description. Existing proposals discussed in this section use systems designed such that a bottleneck transition for transport is suppressed but can be overcome via one of these interactions, this way correlating the particle transport with the transfer of heat, see Fig. 5.
When the system-bath couplings are weak, these systems fall in the category of open quantum systems [234, 235]. A simple description of their dynamics is given in terms of the master equation for the evolution of the system (reduced) density matrix, $\hat{\rho}(t)$. One usually assumes Born-Markov and secular approximations, resulting in the Gorini-Kossakowski-Sudarshan-Lindblad master equation [234, 236]
$\dot{\hat{\rho}}=-\frac{i}{\hbar}[\hat{H}_{S},\hat{\rho}]+\sum_{is}\sum_{qp}W_{qp}^{is}\left(\hat{Y}_{i,qp}\hat{\rho}\hat{Y}_{i,qp}^{\dagger}-\frac{1}{2}\{\hat{Y}_{i,qp}^{\dagger}\hat{Y}_{i,qp},\hat{\rho}\}\right)$ (16)
where the first term in the right hand side represents the coherent dynamics induced by the system Hamiltonian, $\hat{H}_{S}$, and the second term describes the dissipative dynamics induced by the reservoirs, produced by jump operators $\hat{Y}_{i,qp}$ describing transitions $|p\rangle\rightarrow|q\rangle$ due to the interaction $\hat{H}_{SB,i}=\sum_{k_{i}}\lambda_{i}\hat{d}_{\alpha_{i}}^{\dagger}\hat{c}_{k_{i}}+{\rm H.c.}$ with reservoir $i$, where $\lambda_{i}$ is the weak system-bath coupling, $\hat{d}_{\alpha_{i}}^{\dagger}$ creates an electron in quantum dot $\alpha_{i}$ and $\hat{c}_{k_{i}}$ destroys it in reservoir $i$. The transition rates $W_{qp}^{is}$ are given by the Fermi golden rule, where the index $s=\pm$ labels whether the transition involves a particle tunneling into/out of the system [235]:
@@
In the later equation, it is assumed that the energy levels are discrete, which is a good approximation in the weak coupling regime, $W_{qp}^{is}\ll k_{\text{B}}T_{i}/\hbar$. The dc currents are obtained by simply replacing the stationary solution of the master equation, $\dot{\hat{\rho}}=0$, in the above expressions. Regimes with stronger couplings require more involved treatments, e.g., in terms of nonequilibrium Green functions [237, 238].
@@ IV.1 Electron-electron interactions
Capacitively coupled quantum dots ($s$ and $g$) interact via the Hamiltonian $\hat{H}_{U}=U\hat{n}_{s}\hat{n}_{g}$, where $U$ is the interdot Coulomb repulsion and $\hat{n}_{\alpha}=\hat{d}_{\alpha}^{\dagger}\hat{d}_{\alpha}$ is the occupation of dot $\alpha$, such that a series of fluctuations
@@
illustrated in Figs. 5(a) and 5(b), transfers an energy quantum $U$ from $g$ to $s$. This energy is provided by the reservoir coupled to the former dot. When the two system transitions $(0,0)\rightarrow(1,0)$ and $(1,1)\rightarrow(0,1)$ occur through different barriers coupling dot $s$ to terminals $L$ and $R$, the sequence results an electron transported through the conductor, enabling the heat conversion engine [87]. This sequence is relevant when the transport of an electron without interacting with $g$, $(0,0)\rightarrow(1,0)\rightarrow(0,0)$, is suppressed, e.g., if $\Gamma_{0}^{R}\ll\Gamma_{0}^{L}$, with $\Gamma_{n_{g}}^{i}$ being the rate for tunneling events through the barrier $i$ between states $(0,n_{g})$ and $(1,n_{g})$. In this case, the absorption of an energy $U$ activates the transition with a larger rate $\Gamma_{1}^{R}$, see Fig. 5(b). For the process to generate a finite current, the tunneling barriers need to be left-right asymmetric and energy dependent: in the sequential tunneling regime, the current is
@@
Similar processes lead to the mesoscopic Coulomb drag effect in voltage driven sources [239] and thermal drag in temperature biased coupled systems [174]. If the two rates in either of the terms contributing to the current are negligible (e.g., by including additional energy filters [87, 240]), all electrons that absorb an energy $U$ move in the same direction, leading to a Carnot efficient heat engine [87], see also reviews discussing this configuration for more details [7, 8, 241].
In the sequential tunneling regime just discussed, with $\hbar\Gamma_{n}^{i}\ll k_{\text{B}}T_{i}$, phase coherence is lost in every transition. Transport is strongly dominated by nonequilibrium fluctuations, which can be traced in time via charge detectors [242, 243]. This makes not only the particle but also the transferred heat statistics accessible [101]. Time traces of the charge fluctuations reveal the connection of efficiency and crosscorrelations between transported particles and heat quanta transferred from the heat source [121]. It also demonstrates the purely stochastic nature of the dynamics [127], which does not require any moving or internally oscillating parts.
@@ IV.1.1 Quantum correlations
One can also use even-parity superpositions of the occupation of a quantum dot in the proximity of a superconductor to leverage a nonlocal thermoelectric response in the large gap regime, $\Delta\gg k_{\text{B}}T_{i}$. Consider for instance a configuration of coupled quantum dots as sketched in Fig. 5(b) where terminal $R$ is replaced by a large gap superconductor. In this regime, the Coulomb interaction interplays with Andreev reflection processes [247]. The occupation $n_{g}=0,1$ of a capacitively coupled (hot) quantum dot, labelled as $g$, affects not only the energy of the even parity system states $E_{\pm,n_{g}}=A_{\mp}+\varepsilon_{g}$, but also the superposition itself: |±$\left|\pm,n_{g}\right\rangle\propto E_{\pm}\left|0,n_{g}\right\rangle-\Gamma_{S}\left|2,n_{g}\right\rangle$, where $|n_{s},n_{g}\rangle$ are the dot occupation states, $A_{\pm,n_{g}}=\tilde{\varepsilon}_{s}\pm\sqrt{(\tilde{\varepsilon}_{s}+n_{g}U)^{2}+\Gamma_{S}^{2}}$, $\tilde{\varepsilon}_{s}=\varepsilon_{s}+U/2$, $\varepsilon_{\alpha}$ is the energy of dot $\alpha\in\{s,g\}$, and $\Gamma_{S}$ is its coupling to the superconductor. Furthermore, fluctuations in $n_{g}$ induce transitions between them, e.g., $|+,0\rangle\rightarrow|-,1\rangle$ [185]. These properties are sufficient to make the contribution of electrons and holes different in the tunneling evens between the system dot and a normal terminal. As a consequence, a finite current is generated between the normal and the superconducting terminals. Differently from all normal analogues [87, 60], this engine achieves a heat to power conversion without requiring energy dependent tunneling barriers, but rather relying on the stochastic modulation of the system wave function. Additionally, the dependence of |±$|\pm,n_{g}\rangle$ on $n_{g}$ introduces the possibility to modulate the generated current and even its sign by gating the coupled dot. The same effect can also be used as a refrigerator and for hybrid operations [186].
@@
Finally, the nonlocal thermoelectric effect can be used to probe the impact of interactions in the quantum Hall regime. For this, one considers a device with two edge states that are copropagating along a certain distance. With a proper design of quantum point contacts, one can inject electrons from a hot terminal into the outer edge and then separate the two edges such that only the inner edge scatters with a thermoelectric element (e.g., another quantum point contact). If the inner edge connects two terminals at the same temperature, the current between them will be finite only provided some heat has been transferred between the edge channels mediated by the electron-electron interaction of electrons in different channels [95]. Furthermore, one can selectively replace one of the edges at a given point in the interacting region by an equivalent copy, this way avoiding any previously developed crosscorrelations. The generated response is sensitive to this replacement [95], hence the thermoelectric effect allows for the detection of nonequilibrium crosscorrelations of interacting Luttinger liquids.
@@ IV.1.2 Information engines and autonomous Maxwell demons
Interestingly the correlated dynamics of the two coupled subsystems can be interpreted in terms of mutual information [108, 109, 249], triggering connections with information-driven engines [111]. Indeed, focusing on the capacitively coupled quantum dot systems discussed above [87], the two transitions in the source dot, $g$, perform the detection (of the charge in $s$), feedback (the charge in $g$ affecting the tunneling rates in the conductor) and information erasure (restoring the initial state) mechanisms of an autonomous Maxwell demon. This is most clearly evidenced in the equivalent case where the nonlocal thermoelectric response is due to coupling to a colder source [107], where the sequence is reversed: $(n_{1},n_{2})=(0,0)\rightarrow(0,1)\rightarrow(1,1)\rightarrow(1,0)\rightarrow(0,0)$. The cold dot, $g$, is occupied until an electron enters $s$, when it absorbs the energy $U$ to overcome the electrochemical potential $\mu_{g}$ and tunnel out (detection). This changes the rates in the system (backaction) allowing the electron to tunnel out and restores the initial state (reset). In this case, $U$ is transferred from $s$ to $g$, increasing the entropy of the gate.
Differently from what one expects from a Maxwell demon though, a three terminal configuration cannot reduce the entropy of the conductor by, e.g., generating power, without changing its energy: a finite $J_{H}\neq 0$ is necessarily exchanged via the electron-electron interaction in the quantum dots, see Eq. (22). In order to have a conductor subsystem (formed by terminals $L$ and $R$) where the first law is respected while the second is violated), a four terminal configuration is required, with $s$ coupled to two different dots, one hot, the other cold [123, 110, 250, 251]. The thermodynamic discussion is analogous to the one described in Sec. III.2.2 for nonequilibrium states: the entropy reduction in the conductor is compensated by the heat flowing in the coupled terminals. Indeed, if one forces the system quantum dot to reach a thermal state, the demonic effect disappears [123]. However, the role of information is in this case explicit [110, 113].
@@ IV.2 Bosonic excitations
The case where the injected heat comes from a bosonic environment requires a conductor with the same broken asymmetries examined in Sec. IV.1, as discussed for a minimal configuration consisting on a single quantum dot (or molecule) coupled to a phonon bath in Ref. [114]. Most of the subsequent proposals focus in the hopping regime, where the nanostructure is described by a tight binding approach as a series of localized states. The electrons hop from one to another in their way through the conductor. In the case that the nanostructure contains a barrier or a bottleneck transition (e.g., if a site in a one-dimensional chain has a higher energy), transport is suppressed unless additional energy is provided to the electrons from the environment. This environment, that can be treated as a (bosonic) third terminal of the conductor, can take the form of lattice phonons, photons from the electromagnetic fluctuations in nearby conductors or magnons in magnetic insulators, depending on the specific configuration [252].
@@
Most of these works assume sequential transitions where phase coherence is lost after every hop. In this regime, it has been shown that the thermoelectric response is defined by the sites coupling to the electronic terminals [154], which determine how much heat is dissipated in the conductor. This is not expected in a phase coherent conductor where the inner sites affect the molecular superpositions [253]. Respecting the phase coherence at the same time preserving the thermodynamic laws requires a more delicate treatment [256] which may be untreatable as the system size increases.
@@
Similar effects can be due to coupling to fluctuations of the electromagnetic environment (photons) [117, 119, 164] or to thermal photons [258, 259]. This process can be optimized by coupling a double quantum dot to photons in a cavity [118, 260]. The cavity mediated coupling has the practical advantage that the heat bath can be spatially separated from the harvester device, therefore allowing for larger temperature differences between the thermal bath and the conductor. Photons are injected into the cavity by charge fluctuations of an additional double quantum dot coupled to hotter reservoirs. The photon frequency is controlled by the interdot barrier of the hot double quantum dot [118]. Recent experiments along these lines have measured generated powers of $1\,\mathrm{fW}$ with 12% efficiency at electronic temperatures around $40\,\mathrm{mK}$ [261, 66].
Interesting connections can be established with recent proposals of quantum heat engines under continuous monitoring [262, 263, 264], which renders thermodynamic interpretations of quantum measurement processes [265, 266].

### v2

@@ IV Interacting systems
In some quantum systems, single-particle interactions become important and cannot be treated in a mean-field level. This is the case of quantum dot systems weakly coupled to electronic reservoirs [276], where strong Coulomb interactions stabilize the number of allowed electrons (Coulomb blockade regime [277]) up to single-charge fluctuations through the tunnel barriers [278]. For example, Coulomb interactions between electrons in capacitively coupled conductors or interactions with bosonic excitations (phonons, photons, magnons) need to be described microscopically. Existing proposals discussed in this section use systems designed such that a bottleneck transition for transport is suppressed but can be overcome via one of these interactions, this way correlating the particle transport with the transfer of heat quanta, see Fig. 5.
When the system-bath couplings are weak, these systems fall in the category of open quantum systems [279, 280, 281]. A simple description of their dynamics is given in terms of the master equation for the evolution of the system (reduced) density matrix, $\hat{\rho}(t)$. One usually assumes Born-Markov and secular approximations, resulting in the Gorini-Kossakowski-Sudarshan-Lindblad master equation [282, 283, 284]
[…]
where the first term in the right hand side represents the coherent dynamics induced by the system Hamiltonian, $\hat{H}_{S}$, and the second term describes the dissipative dynamics induced by the reservoirs, produced by jump operators $\hat{Y}_{i,qp}$ describing transitions between the system states $|p\rangle\rightarrow|q\rangle$ due to the system-bath interaction $H_{SB}$. For the coupling to the electronic reservoir $i$, this is $\hat{H}_{SB,i}=\sum_{k_{i}}\lambda_{i}\hat{d}_{\alpha_{i}}^{\dagger}\hat{c}_{k_{i}}+{\rm H.c.}$, where $\lambda_{i}$ is the weak system-bath coupling, $\hat{d}_{\alpha_{i}}^{\dagger}$ creates an electron in quantum dot $\alpha_{i}$ and $\hat{c}_{k_{i}}$ destroys it in reservoir $i$. The transition rates $W_{qp}^{is}$ are given by the Fermi golden rule, where the index $s=\pm$ labels whether the transition involves a particle tunneling into/out of the system [280]:
@@
In the later equation, it is assumed that the energy levels are discrete, which is a good approximation in the weak coupling regime, $W_{qp}^{is}\ll k_{\text{B}}T_{i}/\hbar$. The dc currents are obtained by simply replacing the stationary solution of the master equation, $\dot{\hat{\rho}}=0$, in the above expressions. Regimes with stronger couplings require more involved treatments, e.g., via higher order perturbative expansions [285, 286] or in terms of nonequilibrium Green functions [287, 288].
@@ IV.1 Electron-electron interactions
Capacitively coupled quantum dots (one in the two-terminal conductor system, $s$, and the other one, $g$, coupled to a hot terminal, as in Fig. 5(a)) enable the spacial separation of the heat source and the conductor, with the heat exchange being restricted to the mesoscopic region. Electrons in different dots interact via the Hamiltonian $\hat{H}_{U}=U\hat{n}_{s}\hat{n}_{g}$, where $U$ is the interdot Coulomb repulsion and $\hat{n}_{\alpha}=\hat{d}_{\alpha}^{\dagger}\hat{d}_{\alpha}$ is the occupation of dot $\alpha$, such that a series of fluctuations
@@
illustrated in Figs. 5(a) and 5(b), transfers an energy quantum $U$ from $g$ to $s$. This energy is provided by the hot reservoir coupled to the former dot. When the two system transitions $(0,0)\rightarrow(1,0)$ and $(1,1)\rightarrow(0,1)$ occur through different barriers coupling dot $s$ to terminals $L$ and $R$, the sequence results in an electron transported through the conductor, enabling the heat conversion engine [96]. This sequence is relevant when the transport of an electron without interacting with $g$, $(0,0)\rightarrow(1,0)\rightarrow(0,0)$, is suppressed, e.g., if $\Gamma_{0}^{R}\ll\Gamma_{0}^{L}$, with $\Gamma_{n_{g}}^{i}$ being the rate for tunneling events through the barrier $i$ between states $(0,n_{g})$ and $(1,n_{g})$. In this case, the absorption of an energy $U$ activates the transition with a larger rate $\Gamma_{1}^{R}$, see Fig. 5(b). For the process to generate a finite current, the tunneling barriers need to be left-right asymmetric and energy dependent: in the sequential tunneling regime, the current is
@@
This effect has been confirmed experimentally in various settings of coupled quantum dots [68] and open cavities [69, 70]: A small current $\sim$0.2–3 pA is generated which can be reversed by gating the tunneling asymmetry [68, 69] at 0.2–0.8 K temperatures. Similar processes lead to the mesoscopic Coulomb drag effect in voltage driven sources [289, 290, 291, 292, 89, 91] and thermal drag in temperature biased coupled systems [189, 191]. If the two rates in either of the terms contributing to the current are negligible (e.g., by including additional energy filters [96, 293]), all electrons that absorb an energy $U$ move in the same direction, increasing power and leading to a Carnot efficient heat engine [96], see also reviews discussing this configuration for more details [7, 294, 8, 295].
In the sequential tunneling regime just discussed, with $\hbar\Gamma_{n}^{i}\ll k_{\text{B}}T_{i}$, phase coherence is lost in every transition. Transport is strongly dominated by nonequilibrium fluctuations, which can be traced in time via charge detectors [296, 297]. This makes not only the particle but also the transferred heat statistics accessible [112], as recently verified [298]. Time traces of the charge fluctuations reveal the connection of efficiency and crosscorrelations between transported particles and heat quanta transferred from the heat source [133]. It also demonstrates the purely stochastic nature of the dynamics [139], which does not require any moving or internally oscillating parts.
@@ IV.1.1 Quantum correlations
One can also use even-parity superpositions of the occupation of a quantum dot in the proximity of a superconductor to leverage a nonlocal thermoelectric response in the large gap regime, $\Delta\gg k_{\text{B}}T_{i}$. Consider for instance a configuration of coupled quantum dots as sketched in Fig. 5(b) where terminal $R$ is replaced by a large gap superconductor. In this regime, the Coulomb interaction interplays with Andreev reflection processes [302]. The occupation $n_{g}=0,1$ of a capacitively coupled (hot) quantum dot, labelled as $g$, affects not only the energy of the even parity system states $E_{\pm,n_{g}}=A_{\mp}+\varepsilon_{g}$, but also the superposition itself: $\left|\pm,n_{g}\right\rangle\propto E_{\pm}\left|0,n_{g}\right\rangle-\Gamma_{S}\left|2,n_{g}\right\rangle$, where $|n_{s},n_{g}\rangle$ are the dot occupation states, $A_{\pm,n_{g}}=\tilde{\varepsilon}_{s}\pm\sqrt{(\tilde{\varepsilon}_{s}+n_{g}U)^{2}+\Gamma_{S}^{2}}$, $\tilde{\varepsilon}_{s}=\varepsilon_{s}+U/2$, $\varepsilon_{\alpha}$ is the energy of dot $\alpha\in\{s,g\}$, and $\Gamma_{S}$ is its coupling to the superconductor. Furthermore, fluctuations in $n_{g}$ induce transitions between them, e.g., $|+,0\rangle\rightarrow|-,1\rangle$ [201]. These properties are sufficient to make the contribution of electrons and holes different in the tunneling evens between the system dot and a normal terminal. As a consequence, a finite current is generated between the normal and the superconducting terminals. Differently from all normal analogues [96, 68], this engine achieves a heat to power conversion without requiring energy dependent tunneling barriers, but rather relying on the stochastic modulation of the system wave function. Additionally, the dependence of $|\pm,n_{g}\rangle$ on $n_{g}$ introduces the possibility to modulate the generated current and even its sign by gating the coupled dot. The same effect can also be used as a refrigerator and for hybrid operations [202].
@@
Finally, the nonlocal thermoelectric effect can be used to probe the impact of interactions in the quantum Hall regime. For this, one considers a device with two edge states that are copropagating along a certain distance. With a proper design of quantum point contacts, one can inject electrons from a hot terminal into the inner edge and then separate the two edges such that only the outer edge scatters with a thermoelectric element (e.g., another quantum point contact). If the outer edge connects two terminals at the same temperature, the current between them will be finite only provided some heat has been transferred between the edge channels mediated by the electron-electron interaction of electrons in different channels [104]. Furthermore, one can selectively replace one of the edges at a given point in the interacting region by an equivalent copy, this way avoiding any previously developed crosscorrelations. The generated response is sensitive to this replacement [104], hence the thermoelectric effect allows for the detection of nonequilibrium crosscorrelations of interacting Luttinger liquids. Recent experiments confirm this picture [105].
@@ IV.1.2 Information engines and autonomous Maxwell demons
Interestingly the correlated dynamics of the two coupled subsystems can be interpreted in terms of mutual information [119, 303, 120, 304], triggering connections with information-driven engines [122, 305]. Indeed, focusing on the capacitively coupled quantum dot systems discussed above [96], the two transitions in the source dot, $g$, perform the detection (of the charge in $s$), feedback (the charge in $g$ affecting the tunneling rates in the conductor) and information erasure (restoring the initial state) mechanisms of an autonomous Maxwell demon. This is most clearly evidenced in the equivalent case where the nonlocal thermoelectric response is due to coupling to a colder source [118], where the sequence is reversed: $(n_{1},n_{2})=(0,0)\rightarrow(0,1)\rightarrow(1,1)\rightarrow(1,0)\rightarrow(0,0)$. The cold dot, $g$, is occupied until an electron enters $s$, when it absorbs the energy $U$ to overcome the electrochemical potential $\mu_{g}$ and tunnel out (detection). This changes the rates in the system (backaction) allowing the electron to tunnel out and restores the initial state (reset). In this case, $U$ is transferred from $s$ to $g$, increasing the entropy of the gate.
Differently from what one expects from a Maxwell demon though, a three terminal configuration cannot reduce the entropy of the conductor by, e.g., generating power, without changing its energy: a finite $J_{H}\neq 0$ is necessarily exchanged via the electron-electron interaction in the quantum dots, see Eq. (22). References [68] and [306] are then rather interpreted as autonomous information-enabled (demonic) heat engines [122]. In order to have a conductor subsystem (formed by terminals $L$ and $R$) where the first law is respected while the second is violated, a four terminal configuration is required, with $s$ coupled to two different dots, one hot, the other cold [135, 121, 307, 308]. The thermodynamic discussion is analogous to the one described in Sec. III.4 for nonequilibrium states: the entropy reduction in the conductor is compensated by the heat flowing in the coupled terminals. Indeed, if one forces the system quantum dot to reach a thermal state, the demonic effect disappears [135]. However, the role of information is in this case explicit [121, 124].
@@ IV.2 Bosonic excitations
The case where the injected heat comes from a bosonic environment requires a conductor with the same broken symmetries examined in Sec. IV.1, as discussed for a minimal configuration consisting on a single quantum dot (or molecule) coupled to a phonon bath in Ref. [125]. Most of the subsequent proposals focus in the hopping regime, where the nanostructure is described by a tight binding approach as a series of localized states. The electrons hop from one to another in their way through the conductor. In the case that the nanostructure contains a barrier or a bottleneck transition (e.g., if a site in a one-dimensional chain has a higher energy), transport is suppressed unless additional energy is provided to the electrons from the environment. This environment, that can be treated as a (bosonic) third terminal of the conductor, can take the form of lattice phonons, photons from the electromagnetic fluctuations in nearby conductors or magnons in magnetic insulators, depending on the specific configuration [309].
@@
Most of these works assume sequential transitions where the electronic phase coherence is lost after every hop. In this regime, it has been shown that the thermoelectric response is defined by the sites coupling to the electronic terminals [168], which determine how much heat is dissipated in the conductor. This is not expected in a phase coherent conductor where the inner sites affect the molecular superpositions [310]. Respecting the phase coherence at the same time preserving the thermodynamic laws requires a more delicate treatment [313] which may be untreatable as the system size increases.
@@
Similar effects can be due to coupling to fluctuations of the electromagnetic environment (photons) [128, 130, 179] or to thermal photons [315, 316]. This process can be optimized by coupling a double quantum dot to photons in a cavity [129, 317]. The cavity mediated coupling has the practical advantage that the heat bath can be spatially separated from the harvester device over macroscopic distances [318], therefore allowing for larger temperature differences between the thermal bath and the conductor. Photons are injected into the cavity by charge fluctuations of an additional double quantum dot coupled to hotter reservoirs. The photon frequency is controlled by the interdot barrier of the hot double quantum dot [129]. Recent experiments along these lines have measured generated powers of $1\,\mathrm{fW}$ with 12% efficiency at electronic temperatures around $40\,\mathrm{mK}$ [319, 74].
Interesting connections can be established with recent proposals of quantum heat engines powered by continuous monitoring [320, 321, 322], which renders thermodynamic interpretations of the quantum measurement process [323, 324].